A negatively charged particle enters a region between two parallel plates. Each plate (top and bottom) are oppositely charged, hence the exists a uniform electric field between the plates along with a gravitational field). Assume the particle travels the region straight through completely un-deflected and exits. 
a) Determine and identify below) te polarities of the plates and the resulting Electric Field direction inside the plated region needed so the particle travels through un-deflected: 
b) If the particle has a mass of 28.5 mg and a charge of -7.6 uc, determine the Electric Field Strength E needed inside the plated region so the particle travels through un-deflected (Fnet=0 )
Answer #1

a)

here , as there in no deflection ,

the electric force must be in upwards direction to counter the gravity.

as the charge is negative ,

electic field is in downwards direction.

upper plate is positve and lower is negative

b)

Now,

q *E = m*g

7.6 *10^-6 * E = 28.5 *10^-6 * 9.8

E = 36.8 N/C

the eletcic field strength is 36.8 N/C